# RRCANet: Recurrent Reusable-Convolution Attention Network for Infrared Small Target Detection

### v2

@@ IV-A3 False Alarm Rate
${F}_{a}=\frac{Pixel_{false}}{Pixel_{All}}.$ (13)
@@ IV-A4 Receiver Operation Characteristics
Receiver Operation Characteristics (ROC) is used to describe the shifting trends of the detection probability (${P}_{d}$) under varying false alarm rate (${F}_{a}$). ROC indicates the total effect under a sliding threshold.
IV-A5 Parameters
Parameters refers to the total number of learnable parameters in the model, which reflects the model size. Taking the conventional 2D convolution as a example, input feature map is $F\in\mathbb{R}^{C_{in}\times H_{in}\times W_{in}}$, 2D convolution kernel is $C\in\mathbb{R}^{C_{in}\times K_{h}\times K_{w}}$ and output feature map is $O\in\mathbb{R}^{C_{out}\times H_{out}\times W_{out}}$. #$\#Params$ is related to the size of convolution kernel. The calculation formula is
#$\#Params=K_{h}\times K_{w}\times C_{in}\times C_{out},$ (14)
@@ IV-C1 Quantitative Results
For all the compared algorithms, we first obtain their predictions and then suppress low-response noise by setting a threshold to calculate evaluation metrics of detection performance. For traditional methods, the adaptive threshold is calculated according to the approach in [14]. For CNN-based methods, we follow their original papers and adopted their fixed thresholds. We keep all remaining hyper-parameters the same as their original papers.
@@
ResUNet [45] 66.06 90.48 10.63 74.06 96.96 14.97 80.96 98.06 4.377 0.9141
AMFU-net [46] 66.70 92.86 16.02 74.26 97.34 14.40 80.33 97.89 4.132 0.4730
DNANet [14] 67.70 91.58 18.69 76.63 95.44 26.55 81.45 97.28 3.252 4.697
ResUNet with RuCB 67.81 92.86 9.489 75.13 95.82 8.842 83.12 98.58 1.096 0.979
AMFU with RuCB 66.19 92.59 5.436 74.30 95.06 5.419 81.48 97.77 1.796 0.489
DNANet with RuCB 69.29 91.50 5.162 75.84 96.58 5.134 80.80 98.12 2.394 4.697
@@
As for computational efficiency, RRCA-Net with ResNet-10 only has 0.103M parameters and still achieves a advanced detection performance under such low computational cost. That is because that reusing convolution does not introduce extra parameters, and can also effectively extract deep features. It helps to extract informative multi-level features after multiple feature refinement and thus introduces accurate detection.
Overall, our proposed RRCA-Net achieves a superior balance between detection performance and computational efficiency.
@@ IV-D Improvements on Existing Methods
In Table II, we insert our proposed RuCB to three popular IRSTD methods (e.g., ResUNet, AMFU-net and DNANet). After using the module, the detection performance of networks improves (i.e., IoU and $P_{d}$ increase, $F_{a}$ decrease). Especially, the ability to suppress false alarm is greatly improved. That is because that multiple iterations allow the networks to further refine the features and eliminate the interfering information. At the same time, model size almost remains unchanged. Experiments can also verify that our strategy can be transferred to CNN-based methods.
@@ IV-E1 The Backbone
IV-E2 Channel and Depth
Model Channel///Depth Performance metrics ##\##Params FLOPs
$IoU$ $P_{d}$ $F_{a}$
C=4,8,16,32;$l$=4 73.06 97.34 24.39 0.052 0.28
C=16,32,64,128;$l$=4 74.76 95.82 21.61 0.799 4.00
C=8,16,32,64;$l$=4 75.41 98.10 7.558 0.202 1.04
C=8,16,32;$l$=3 74.11 96.96 13.98 0.078 1.02
C=8,16,32,64,128;$l$=5 75.39 96.96 20.82 0.785 1.05
C=4,8,16,32,64;$l$=5 72.60 98.10 23.39 0.200 0.28
C=16,32,64,128,256;$l$=5 75.64 96.96 22.82 3.122 4.06
[…]
@@ IV-E2 Channel and Depth
Number of iterations Evaluation Metircs ##\##Params FLOPs
@@
As discussed before, C=8,16,32,64 and $l$=4 is the most appropriate channel and depth configurations, respectively.
IV-E3 The Reusable Convolution Block (RuCB)
The reusable convolution block is used to employs recurrent convolution to iteratively refine the feature maps extracted at different times while no extra parameters are introduced. This block helps to maintain fine-grained features at the finest scale level while eliminate superfluous information. To verify the effectiveness of RuCB, we train and test RRCA-Net with different iterative times $n$.
@@ IV-E3 The Reusable Convolution Block (RuCB)
TABLE V: Results achieved by variants of RRCA-Net and FSM on the NUAA-SIRST dataset. $DS$ means the deep supervision.
Model Evaluation Metircs ##\##Params FLOPs
@@
Table IV shows the results achieved by RRCA-Net with changed recurrent times. When $n=0$, network does not reuse convolution kernels, it is the encoder of the plain U-shape network. We can see that segmentation maps is iterated once in the loop, IoU value increases, and false alarm rate decreases. The best performance metrics are obtained at $n=2$, where the $IoU$, $P_{d}$, and $F_{a}$ values are improved by $1.87\%$, $1.14\%$ and get decreases of $9.202\times 10^{-6}$ compared to RRCA-Net without RuCB. The reason is that RuCB repeatedly refines features in recurrent iterations to maintain the completeness and comprehensiveness of target information for more advanced performance. However, excessive iteration number (i.e., N $\geq$ 3) will make the RRCA-Net suffers from performance degradation, even worse than network without RuCB. We think that too many reusing times will not improve the detection ability, but weaken the network’s ability to suppress complex background.
@@
##\##Params in Table IV have a slight rise with the increase of iterations. That is because that RuCB only reuses the convolution kernels, and new learnable parameters are still added in Batch Normalization layers when iterations occur. And point-wise convolutions in DIAAM also introduce a slight increase in parameters. Each iteration increases ##\##Params by 9K, while FLOPs unsurprisingly increases by 0.11G. Both increase are acceptable.
Figure 10: Two variants of DIAAM. (a) RRCA-Net w/o DIAAM©. (b) RRCA-Net w/o DIAAMdirect-$\oplus$. Orange represents bilinear upsample and concatenate operations, purple represents bilinear upsample, $1\times 1$ convolution and summation operations.
TABLE VI: Results achieved by main variants of RRCA-Net and DIAAM on the NUAA-SIRST datasets. © and $\oplus$ respectively mean concatenate and element-wise summation as feature fusion methods. $CA$ means the channel attention mechanism.
Model Evaluation Metircs ##\##Params FLOPs
IoU $P_{d}$ $F_{a}$
RRCA-Net w/o DIAAM© 74.11 96.58 25.669 0.199 1.60
RRCA-Net w/o DIAAMdirect-$\oplus$ 73.26 96.58 26.311 0.190 0.98
RRCA-Net w/o CA 73.17 96.20 20.892 0.201 1.04
RRCA-Net 75.41 98.10 7.558 0.202 1.04
@@ IV-E5 The Dual Interactive Attention Aggregation Module (DIAAM)
  • RRCA-Net w/o channel attention: We remove the channel attention operation in high-level branch of DIAAM to analyze its contribution.
As shown in Table VI, RRCA-Net w/o DIAAMdirect-$\oplus$ suffers decreases of 2.15%, 1.52%, and a increase of 18.753 ×$\times 10^{-6}$ in terms of $IoU$, ${P}_{d}$, and ${F}_{a}$ on the NUAA-SIRST dataset. Similar results can also be observed for RRCA-Net w/o DIAAM©. The reason is that DIAAM gradually aggregates the upper-layer and lower-layer features to maintain the elaborate information such as centroid location and shape contour. Conversely, upsampling is prone to generate deviations.
[…]
@@ IV-F Benefits of DpT-k Loss
In Table VIII, we can find that single loss function performs poor on there metrics, while Dice$+$Top-k loss also achieves the highest IoU value close to DpT-k loss. But $P_{d}$ and $F_{a}$ achieved by Dice$+$Top-k loss are worse. DpT-k loss achieves the advanced detection performance, especially obtaining a low false alarm. As shown in Fig. 12, RRCA-Net trained with DpT-k loss has a strong ability to locate dim targets and suppress the false alarms. That is because that DpT-k loss can combine the advantages of Dice loss, Poly loss and Top-k loss, the approach allows loss function to help RRCA-Net focuses on hard samples and multi-scale targets. Meanwhile the combination of shape constraints at physical (i.e., the contribution of Dice) and distribution constraints at the mathematical level (i.e. the contribution of Poly and Top-k) guide the network to converge more stably and get better performance.
@@ V CONCLUSION
In this paper, we propose a RRCA-Net to achieve high-performance IRST detection with low computation cost. Different from previous lightweight methods, we are inspired by the reusing mechanism and designed a reusable network with cascaded attention module to achieve high-level feature refinement and enhancement, respectively. Moreover, we develop a target characteristic-inspired loss function to guide the convergence toward precise location and contour description. Experiments on three datasets have shown that RRCA-Net achieves comparable performance to the state-of-the-art methods, and introduces consistent performance improvement for several popular IRSTD methods.

### v3

@@ IV-A3 False Alarm Rate
${F}_{a}=\frac{Pixel_{false}}{Pixel_{All}},$ (13)
@@ IV-A4 Receiver Operation Characteristics
Receiver Operation Characteristics (ROC) is used to describe the shifting trends of the detection probability (${P}_{d}$) under varying false alarm rate (${F}_{a}$). ROC indicates the total effect under a sliding threshold, comprehensively reflecting the overall performance of the network on ${P}_{d}$ and ${F}_{a}$.
IV-A5 #\#Params
Parameters refers to the total number of learnable parameters in the model, which reflects the model size. Taking the conventional 2D convolution as an example, input feature map is $F\in\mathbb{R}^{C_{in}\times H_{in}\times W_{in}}$, 2D convolution kernel is $C\in\mathbb{R}^{C_{in}\times K_{h}\times K_{w}}$ and output feature map is $O\in\mathbb{R}^{C_{out}\times H_{out}\times W_{out}}$. #Params\#Params is related to the size of convolution kernel. The calculation formula is
$Params=K_{h}\times K_{w}\times C_{in}\times C_{out},$ (14)
@@
IV-A6 FLOPs
FLOPs refer to the total number of float point operations performed during the model’s execution, including addition, subtraction, multiplication and division operations. Taking the conventional 2D convolution as an example, FLOPs are related to the input feature and convolution kernels. The calculation formula is
$FLOPs=2\times K_{h}\times K_{w}\times C_{in}\times C_{out}\times H_{in}\times W_{in}.$ (15)
@@ IV-C1 Quantitative Results
For all the compared algorithms, we first obtain their predictions and then suppress low-response noise by setting a threshold to calculate evaluation metrics of detection performance. Traditional methods involve calculating the corresponding adaptive threshold. For CNN-based methods, we follow their original papers and adopted their fixed thresholds. We keep all remaining hyper-parameters the same as their original papers.
@@
ResUNet [52] 66.06 90.48 10.63 74.06 96.96 14.97 80.96 98.06 4.377 0.9141 1.27
AMFU-net [53] 66.70 92.86 16.02 74.26 97.34 14.40 80.33 97.89 4.132 0.4730 5.09
DNANet [14] 67.70 91.58 18.69 76.63 95.44 26.55 81.45 97.28 3.252 4.697 14.05
ResUNet with RuCB 67.81 92.86 9.489 75.13 95.82 8.842 83.12 98.58 1.096 0.979 1.68
AMFU with RuCB 66.19 92.59 5.436 74.30 95.06 5.419 81.48 97.77 1.796 0.489 5.77
DNANet with RuCB 69.29 91.50 5.162 75.84 96.58 5.134 80.80 98.12 2.394 4.697 15.04
@@
As for computational efficiency, RRCA-Net with ResNet-10 only has 0.103M parameters and still achieves a advanced detection performance under such low model parameters. That is because that reusing convolution does not introduce extra parameters, and can also effectively extract deep features. It helps to extract informative multi-level features after multiple feature refinement and thus introduces accurate detection. But FLOPs increase because of using multiple reusing to process features.
Overall, our proposed RRCA-Net achieves a balance between detection performance and model size.
@@ IV-D Improvements on Existing Methods
In Table II, we insert our proposed RuCB to three popular IRSTD methods (e.g., ResUNet, AMFU-net and DNANet). After using the module, the detection performance of networks improves (i.e., IoU and $P_{d}$ increase, $F_{a}$ decrease). Especially, the ability to suppress false alarm is greatly improved. That is because that multiple iterations allow the networks to further refine the features and eliminate the interfering information. At the same time, the number of parameters of the network almost remains unchanged. The multiple iterations result in more floating point operations, but the increase in FLOPs is acceptable. Experiments can also verify that our strategy can be transferred to various CNN-based methods and improve detection performance under low computational cost.
@@ IV-E Ablation Study
Model Channel//Depth Performance metrics #\#Params FLOPs
IoUIoU $P_{d}$ $F_{a}$
C=4,8,16,32;ll=4 73.06 97.34 24.39 0.052 0.28
C=16,32,64,128;ll=4 74.76 95.82 21.61 0.799 4.00
C=8,16,32,64;ll=4 75.41 98.10 7.558 0.202 1.04
C=8,16,32;ll=3 74.11 96.96 13.98 0.078 1.02
C=8,16,32,64,128;ll=5 75.39 96.96 20.82 0.785 1.05
C=4,8,16,32,64;ll=5 72.60 98.10 23.39 0.200 0.28
C=16,32,64,128,256;ll=5 75.64 96.96 22.82 3.122 4.06
@@ IV-E1 The Backbone
Number of iterations Evaluation Metircs #\#Params FLOPs
@@ IV-E2 Channel and Depth
As discussed before, C=8,16,32,64 and ll=4 is the most appropriate channel and depth configurations, respectively.
@@
TABLE V: Results achieved by variants of RRCA-Net and FSM on the NUAA-SIRST dataset. DSDS means the deep supervision.
Model Evaluation Metircs #\#Params FLOPs
@@
IV-E3 The Reusable Convolution Block (RuCB)
Reusable operation allows the same set of convolution kernels to process feature maps multiple times, achieving progressive depth refinement of detailed features. This is equivalent to functional network deepening, without introducing extra convolution kernel parameters. The reusable convolution block is an important component for achieving recurrent reuse. This block helps to maintain fine-grained features at the finest scale level while eliminate superfluous information. To verify the effectiveness of RuCB, we train and test RRCA-Net with different iterative times $n$.
Table IV shows the results achieved by RRCA-Net with changed recurrent times. When $n=0$, network does not reuse convolution kernels, it is the encoder of the plain U-shape network. We can see that segmentation maps is iterated once in the loop, IoU value increases, and false alarm rate decreases. The best performance metrics are obtained at $n=2$, where the $IoU$, $P_{d}$, and $F_{a}$ values are improved by $87\%, 1.1$$14\%$ and get decreases of $9.202\times 10^{-6}$ compared to RRCA-Net without RuCB. The reason is that RuCB repeatedly refines features in recurrent iterations to maintain the completeness and comprehensiveness of target information for more advanced performance. However, excessive iteration number (i.e., N $\geq$ 3) will make the RRCA-Net suffers from performance degradation, even worse than network without RuCB. We think that too many reusing times will not improve the detection ability, but weaken the network’s ability to suppress complex background.
@@ IV-E3 The Reusable Convolution Block (RuCB)
#\#Params in Table IV have a slight rise with the increase of iterations. That is because that RuCB only reuses the convolution kernels, and new learnable parameters are still added in Batch Normalization layers when iterations occur. And point-wise convolutions in DIAAM also introduce a slight increase in parameters. Each iteration increases #\#Params by 9K, while FLOPs unsurprisingly increases by 0.11G. Both increase are acceptable.
[…]
@@ IV-E5 The Dual Interactive Attention Aggregation Module (DIAAM)
  • RRCA-Net w/o CA: We remove the channel attention operation in high-level branch of DIAAM to analyze its contribution.
  • RRCA-Net with CA4: We change the scaling ratio of the shared MLP in channel attention. The ratio factor is set to 4. The original factor of RRCA-Net is 8.
  • RRCA-Net with re-weight: We change the fusion strategy between two branches. Specifically, generating wights from high-level features and then weighting the shallow-level features precessed by channel attention.
TABLE VI: Results achieved by main variants of RRCA-Net and DIAAM on the NUAA-SIRST datasets. © and $\oplus$ respectively mean concatenate and element-wise summation as feature fusion methods. CACA means the channel attention mechanism.
Model Evaluation Metircs #\#Params FLOPs
IoU $P_{d}$ $F_{a}$
RRCA-Net w/o DIAAM© 74.11 96.58 25.669 0.199 1.60
[…]
RRCA-Net w/o CA 73.17 96.20 20.892 0.201 1.04
RRCA-Net with CA4 75.17 96.20 18.967 0.202 1.04
RRCA-Net with re-weight 73.29 96.96 24.100 0.202 1.04
RRCA-Net 75.41 98.10 7.558 0.202 1.04
[…]
@@ IV-F Benefits of DpT-k Loss
In Table VIII, we can find that the single losses (e.g., Dice Loss[44], SoftIoU Loss[45]) perform worse than the joint losses on three metrics. As to joint losses, Dice++Focal loss [46] and FocalIoU loss [21] show the conventional performance and does not reach the optimum. Dice++Top-k loss also achieves the highest IoU value close to DpT-k loss. But $P_{d}$ and $F_{a}$ achieved by Dice++Top-k loss are worse. DpT-k loss achieves the advanced detection performance, especially obtaining a low false alarm. As shown in Fig. 12, RRCA-Net trained with DpT-k loss has a strong ability to locate dim targets and suppress the false alarms. That is because that DpT-k loss can combine the advantages of Dice loss, Poly loss and Top-k loss, the approach allows loss function to help RRCA-Net focuses on hard samples and multi-scale targets. Meanwhile the combination of shape constraints at physical (i.e., the contribution of Dice) and distribution constraints at the mathematical level (i.e. the contribution of Poly and Top-k) guide the network to converge more stably and get better performance.
@@
Dice Loss + Focal Loss 74.13 96.58 21.034
FocalIoU Loss 72.28 95.82 11.765
@@ V CONCLUSION
In this paper, we propose a RRCA-Net to achieve high-performance IRST detection with low computation cost. Different from previous lightweight methods, we are inspired by the reusing mechanism and designed a reusable network with cascaded attention module to achieve high-level feature refinement and enhancement, respectively. Moreover, we develop a target characteristic-inspired loss function to guide the convergence toward precise location and contour description. Experiments on three datasets have shown that RRCA-Net achieves comparable performance to the state-of-the-art methods, and introduces consistent performance improvement for several popular IRSTD methods. However, an obvious limitation of recurrent reuse is that FLOPs will increase. In the future, we will consider reusing the encoder as a decoder to further reduce the model size. And we attempt to introduce reuse operations into the transformer structure.